OpenStudy (anonymous):

What is the solution? -6x + 4 =-8x - 12

OpenStudy (austinl):

We have, \(-6x+4=-8x-12\) We need to combine like terms, \(-6x+8x=-12-4\) Can you take it from there?

OpenStudy (anonymous):

i got -6

OpenStudy (austinl):

\((-6+8)x=-16\) What is -6+8?

OpenStudy (pratyush5):

do it just like the previous question. keep the variables on one side and the constants on other side